OpenStudy (anonymous):

4. Given the linear equation y = -6x – 2, find the y-coordinates of the points (-1, ), (-2, ), and (1, ). Please show all of your work. Plot these points and graph the linear equation.

OpenStudy (anonymous):

put instead of x the respective values and calculate y for x=-1 y=-6(-1)-2=6-2=4 can you do the rest?